OpenStudy (anonymous):

For a positive integer n, if both n and n + 2 are prime, then they are known as twin primes. For example, 59 and 61 are twin primes. Whether or not there are an infinite number of twin primes is a famous unsolved problem in number theory. Find all positive integers n such that n, n + 2, and n + 4 are all prime.

OpenStudy (anonymous):

Hint: Try looking at them mod 3.

OpenStudy (anonymous):

well it follows that \(n,n+2,n+4\) must be a sequence of odd numbers... one such (immediate) pair is given by \(1,3,5\), another is \(3,5,7\), ...

OpenStudy (anonymous):

If n is a prime congruent to 1 mod 3, will n+2 and n+4 still be prime? What about if n is congruent to 2?

OpenStudy (anonymous):

oops \(1\) is not prime...

OpenStudy (anonymous):

@joemath314159 is correct... considering \(\mod3\) is natural because the range of our values is \(4\)

OpenStudy (zzr0ck3r):

@joemath314159 is always correct:)

OpenStudy (anonymous):

Let n be a prime such that \(n\equiv 1\pmod{3}\) Then \(n+2\equiv 1+2\equiv 0\pmod{3}\) and we see that n+2 isn't prime (its divisible by 3). So n cant be equivalent to 1 modulo 3. Now we try \(n\equiv 2\pmod{3}\). It follows that \(n+4\equiv 2+4\equiv 0\pmod{3}\), so n+4 isnt prime. So n can't be equivalent to 2 modulo 3. That leaves \(n\equiv 0\pmod{3}\), but the only prime that is equivalent to 0 modulo 3 is 3 itself. Hence if n, n+2, and n+4 are all primes, then n must equal 3. This gives 3, 5, and 7 as the only primes that satisfy the condition.

OpenStudy (anonymous):

consider that \(n\equiv0,1,2\mod 3\) hence \(n=3k,3k+1,\text{ or }3k+2\). if \(n=3k\) then \(n\) is not prime unless \(k=1\). if \(n=3k+1\) then \(n+2=3k+3=3(k+1)\) is not prime unless \(k+1=1\). if \(n=3k+2\) then \(n+4=3k+6=3(k+2)\mod 3\)

OpenStudy (anonymous):

damnit @joemath314159 i wrote that out earlier but I went out to help my dad

OpenStudy (anonymous):

lolol, my bad >.< I'm a little busy too, im packing.

OpenStudy (zzr0ck3r):

ill medal the guy who doesn't get metaled

OpenStudy (anonymous):

We metal'd each other :P

OpenStudy (zzr0ck3r):

thats love

OpenStudy (anonymous):

and respect :) oldrin's answers are spot on and interesting.

OpenStudy (anonymous):

unless \(k+2=1\)... and then it remains to be shown that \(k=0\) gives prime \(3(k+1)\) but nonprime \(3k\) and similarly \(k=-1\) gives prime \(3(k+2)\) but nonprime \(3k\)... it must be that \(k=1\) hence \(n=3,n+2=5,n+4=7\)

OpenStudy (anonymous):

hehe I try

OpenStudy (anonymous):

I really should log off thought lol, I gotta finish packing >.<

OpenStudy (zzr0ck3r):

yeah , I like watching you two work.

OpenStudy (zzr0ck3r):

when do you leave joe?

OpenStudy (anonymous):

Monday o.O

OpenStudy (zzr0ck3r):

o wow, and what part of texas?

OpenStudy (zzr0ck3r):

its gonna be soooooooooooooooooooooooooooooooo hot

OpenStudy (anonymous):

I'm moving to College Station. Inbetween Austin in Houston

OpenStudy (zzr0ck3r):

nice, I was born in tyler......so hot

OpenStudy (anonymous):

Yeah, its 103 here in San Antonio right now >.<

OpenStudy (anonymous):

ouch :(

OpenStudy (zzr0ck3r):

ahh I didnt realize you were in texas now

OpenStudy (zzr0ck3r):

well at least log on and let us know you got there safely.

OpenStudy (anonymous):

Sure thing :) I'm out, see you all later.

OpenStudy (anonymous):

jeez and I thought Florida was too warm

OpenStudy (anonymous):

Augghhh!!! Who to give the medal to?

OpenStudy (anonymous):

@joemath314159

OpenStudy (zzr0ck3r):

i lived in florida al well. its got nothing on texas imo

OpenStudy (anonymous):

imo, Florida is worse, because its super humid. It might not get as hot, but at least its dry hot here lolol.

OpenStudy (zzr0ck3r):

yeah true

OpenStudy (anonymous):

and joemath is going to say oldrin.bataku

OpenStudy (anonymous):

this is my life :)

OpenStudy (zzr0ck3r):

@orple8 pick one, and ill give the other

OpenStudy (anonymous):

Thanks :)

OpenStudy (zzr0ck3r):

i dont think either of them care anyway:)

OpenStudy (anonymous):

yeah I mostly just enjoy doing the problems

OpenStudy (anonymous):

true. :)